The current risk-free rate is 4.0% and the market risk premium is 5.0%. If Ford Motor Company common stock's beta coefficient is estimated at 1.2, then what is the required rate of return for Ford's stock?

A 9.0%
B 5.0%
C 10.0%
D 4.0%
Answer #1

Calculation of required rate of return:

Required rate of return= risk free rate+ beta*market risk premium

Required return= 4+1.2*5= 4+6= 10

So correct answer is C) 10.0%